# PolyNet: Learning Diverse Solution Strategies for Neural Combinatorial Optimization

### v1

@@ 4.1 Problems
The CVRPTW is an extension of the CVRP that introduces temporal constraints limiting when a customer can receive deliveries from a vehicle. Each customer $i$ is associated with a time window, comprising an earliest arrival time $e_{i}$ and a latest allowable arrival time $l_{i}$. Vehicles may arrive early at a customer $i$, but they must wait until the specified earliest arrival time $e_{i}$ before making a delivery. The travel duration between customer $i$ and customer $j$ is calculated as the Euclidean distance between their locations, and each delivery has a fixed duration. All vehicles initiate their routes from the depot at time 00. We use the CVRP instance generator outlined in Queiroga et al. (2022) to generate customer positions and demands, and we adhere to the methodology established in Solomon (1987) for generating the time windows. It’s essential to note that customer positions are not sampled from a uniform distribution; instead, they are clustered to replicate real-world scenarios. Further details for instance generation are given in Appendix F.
@@
4.2 Solution Diversity During Training
To assess the effectiveness of our diversity mechanism, we conduct short training runs of PolyNet for the TSP, CVRP, and CVRPTW with varying values of $K$. As a baseline we report results for the training of the POMO model. For all runs, we use a batch size of $480$ and a learning rate set at $10^{-4}$. To ensure a stable initial state for training runs, we start all runs from a trained POMO model. For PolyNet, we initialize the PolyNet layer weights to zero, minimizing initial randomness. During training, we perform regular evaluations on a separate validation set of $10,000$ instances, sampling $800$ solutions per instance.
Figure 3: Validation performance during training (log scale).
@@ 4.3 Search Performance
We conduct an extensive evaluation of PolyNet’s search performance, benchmarking it against state-of-the-art neural CO methods. For the considered routing problems, we train separate models for problem instances of size $100$ and $200$, and then evaluate the models trained on $n$ = $100$ using instances with $100$ and $150$ nodes, and the models trained on $n$ = $200$ using instances with $200$ and $300$ nodes. We can thus assess the model’s capability to generalize to instances that diverge from the training data. Throughout our evaluation, we employ the instance augmentation technique introduced in Kwon et al. (2020).
For the training of PolyNet models, we set the parameter $K$ to $128$ across all problems. We use a learning rate of $10^{-4}$ for the TSP and CVRP and $10^{-5}$ for the CVRPTW. For instances of size $n$$=$$100$, we train our models for $300$ epochs ($200$ for the TSP), with each epoch comprising $4\times 10^{8}$ solution rollouts. Note that we warm-start the training for these models using previously trained POMO models to reduce training times (results for cold-starting training can be found in Appendix C). For instances with $n$$=$$200$, we start training based on the $n$ = $100$ PolyNet models, running $40$ additional training epochs ($20$ for the TSP). To optimize GPU memory utilization, we adjust the batch size separately for each problem and its dimensions.
@@
We use the 10,000 test instances with $n$ = $100$ from Kool et al. (2019) and test sets consisting of $1,000$ instances from Hottung et al. (2021) for $n$ = $150$ and $n$ = $200$. For $n$ = $300$, we generate new instances. As a baseline, we use the exact solver Concorde (Applegate et al., 2006) and the heuristic solver LKH3 (Helsgaun, 2017). Additionally, we also compare to DPDP (Kool et al., 2022b) and the diversity-focused method MDAM (Xin et al., 2021) with a beam search width of 256.
@@
Method Test ($10$K instances) Test ($1$K instances) Generalization ($1$K instances)
@@
LKH3333 7.765 0.000% 8h 10.687 0.000% 3h 9.346 0.000% 99m 12.949 0.000% 5h
Unguided POMO  greedy 7.775 0.135% 1m 10.770 0.780% $<$1m 9.393 0.494% $<$1m 13.216 2.061% 1m
@@
CVRP
Similar to the TSP, we use the test sets from Kool et al. (2019) and Hottung et al. (2021). As a baseline, we use LKH3 (Helsgaun, 2017) and the state-of-the-art (OR) solver HGS (Vidal et al., 2012; Vidal, 2022). We compare to the same learning methods as for the TSP with the addition of DACT (Ma et al., 2021).
@@
Method Test ($10$K instances) Test ($1$K instances) Generalization ($1$K instances)
@@
LKH3333 15.646 0.53% 6d 22.003 1.09% 25h 19.222 0.88% 20h 28.157 1.51% 34h
Unguided POMO  greedy 15.754 1.23% 1m 22.194 1.97% $<$1m 19.684 3.30% $<$1m 28.627 3.21% 1m
@@
CVRPTW
For the CVRPTW, we use the state-of-the-art CO solver PyVRP (Wouda et al., 2024) (version 0.5.0) as a baseline stopping the search after 1,000 iterations without improvement. We compare to a POMO implementation that we adjusted to solve the CVRPTW by extending the node features with the time windows and the context information used at each decoding step by the current time point. These models are trained for 50,000 epochs, mirroring the training setup used for the CVRP.
@@
Method Test ($10$K instances) Test ($1$K instances) Generalization ($1$K instances)
@@
Unguided POMO greedy 13,120 4.67% 1m 19,656 6.70% 1m 18,670 7.25% $<$1m 28,022 8.90% 2m
@@
4.4 Search Diversity
To evaluate the diversity mechanism of PolyNet, we compare the diversity of solutions generated by PolyNet to those generated by POMO. More specifically, for a subset of test instances, we sample $100$ solutions per instance with both approaches and compare the solutions found on the basis of their uniqueness and their cost. We calculate the uniqueness of a solution by using the average of the broken pairs distance (Prins, 2009) to all other 99 generated solutions.
Figure 4 shows the results for the first three CVRP instances from the test set. Ideally, we seek solutions with low costs and high uniqueness (i.e., solutions positioned in the top-left corner of the plot). Notably, the majority of solutions laying on the Pareto front of these two objectives originate from PolyNet. In summary, PolyNet is consistently able to generate solutions with higher uniqueness and lower costs, demonstrating its ability to learn strategies that generate not only diverse solutions but also high-quality solutions. See Appendix D for additional visualizations of all three routing problems.
Figure 4: Solution diversity vs. costs for three CVRP instances.
4.5 Ablation Study: Forcing the First Move
In contrast to most state-of-the-art methods, PolyNet does not force diverse first construction actions to enhance exploration during training and testing. Instead, PolyNet relies solely on its built-in diversity mechanism, which does not assume symmetries in the solution space. To assess the efficacy of this approach, we compare the performance of PolyNet with and without forced first move selection when sampling $64\times 100$ solutions per instance.
Table 4 shows the results for all problems with $n$ = $100$. Remarkably, across all scenarios, allowing PolyNet to select the first move yields superior performance compared to forcing the first move. This finding is particularly striking for the TSP, where the first move does not affect solution quality. Furthermore, PolyNet with forced first move selection outperforms Poppy (which also enforces the first move), underscoring that PolyNet’s single-decoder architecture delivers better results than Poppy’s multi-decoder approach.
@@ 4.5 Ablation Study: Forcing the First Move
Obj. Gap Time Obj. Gap Time Obj. Gap Time
PolyNet    Free first move 7.765 0.000% 4m 15.640 0.49% 5m 12,876 2.73% 5m
              Forced first move 7.765 0.006% 4m 15.655 0.59% 5m 12,909 3.00% 6m
Poppy 7.766 0.015% 4m 15.685 0.78% 5m 12,969 3.47% 5m
@@ 5 Conclusion
Regarding our approach’s limitations, we acknowledge that the computational complexity of the attention mechanism we employ restricts its applicability to instances with less than 1,000 nodes. However, it is essential to emphasize that the problem sizes examined in this paper for the CVRP(TW) remain challenging for traditional CO solvers and are highly significant in real-world applications. Furthermore, we note that the black-box nature of the PolyNet’s decision-making may be unacceptable in certain decision contexts.
Acknowledgments
This work is partially supported by Deutsche Forschungsgemeinschaft (DFG) grant 521243122. Additionally, the authors gratefully acknowledge the funding of this project by computing time provided by the Paderborn Center for Parallel Computing (PC2). Furthermore, some computational experiments in this work have been performed using the Bielefeld GPU Cluster. We thank the Bielefeld HPC.NRW team for their support.
@@ Appendix A Experimental Results for the FFSP
To train the PolyNet models, we use a learning rate set to $10^{-4}$, and batch sizes of $256$, $128$, and $32$ for instances comprising $n=20$, $50$, and $100$ jobs, respectively. We set the parameter $K$ to $24$ across all problem sizes. The training runs start from trained MatNet models and continue for $100$ ($20$ for $n=100$) epochs, with each epoch comprising $24\times 10^{5}$ solution rollouts.
@@
For the FFSP, we use the randomly generated test sets from Kwon et al. (2021) and comparison algorithms. In addition to MatNet (Kwon et al., 2021), we compare our method to CPLEX CPLEX-Optimization-Studio (2020) with mixed-integer programming models and metaheuristic solvers. Table 5 provides the results for the FFSP. The gaps are reported with respect to PolyNet with ×$\times 128$ augmentation result, as optimal solutions are not available. The results demonstrate PolyNet’s consistent improvement in performance across all problem sizes. Notably, PolyNet outperforms MatNet-based solvers for the FFSP on all problem sizes and produces solutions of even higher quality within similar run times. Given the different architecture of MatNet compared to those used for routing problems, the results highlight that the performance gains from PolyNet are not confined to a specific network architecture. Moreover, these gains extend beyond routing problems, indicating PolyNet’s broad applicability across various CO problems.
@@ Appendix B Ablation Study: PolyNet Layers
Appendix D Additional Solution Diversity vs. Costs Plots
We provide additional plots that show the diversity and costs of solutions sampled with PolyNet and POMO for the TSP, CVRP, and CVRPTW in Figures 7-9. For each problem we report results for the first six test set instances. The results further reinforce the notion that PolyNet generates solutions with higher diversity and lower costs than POMO.
@@
Figure 8: Solution diversity vs. costs for the CVRP.
@@
Figure 9: Solution diversity vs. costs for the CVRPTW.
@@ Appendix F CVRPTW Instance Generation
Customer positions are generated using the clustered setting (configuration 2) and customer demands are based on the small, large variance setting (configuration 2). The depot is always centered (configuration 2). It is worth noting that the instance generator samples customer positions within the 2D space defined by $[0,999]^{2}$. Independently from the instance generator, vehicle capacities are set at $50$ for instances involving fewer than $200$ customers and increased to $70$ for instances with $200$ or more customers.
To generate the time windows, we adhere to the procedure outlined by Solomon (1987) for instances with randomly clustered customers (i.e., we do not utilize the 3-opt technique to create reference routes). We randomly generate time windows $(e_{i},l_{i})$ for all customers, and set $2400$ as the latest possible time for a vehicle to return to the depot. The time window generation process, as described in Solomon (1987), limits the time windows to ensure feasibility (e.g., by selecting $l_{i}$ so that there is always sufficient time for servicing the customer and returning to the depot). The center of the time window is uniformly sampled from range defined by these limits. We set the maximum width of the time window to $500$ and the service duration to $50$. These parameter values have been deliberately chosen to strike a balance between the constraints of vehicle capacity and time windows, requiring both aspects to be considered during the solution generation process.

### v2

@@ 4.1 Problems
The CVRPTW is an extension of the CVRP that introduces temporal constraints limiting when a customer can receive deliveries from a vehicle. Each customer $i$ is associated with a time window, comprising an earliest arrival time $e_{i}$ and a latest allowable arrival time $l_{i}$. Vehicles may arrive early at a customer $i$, but they must wait until the specified earliest arrival time $e_{i}$ before making a delivery. The travel duration between customer $i$ and customer $j$ is calculated as the Euclidean distance between their locations, and each delivery has a fixed duration. All vehicles initiate their routes from the depot at time 0. We use the CVRP instance generator outlined in Queiroga et al. (2022) to generate customer positions and demands, and we adhere to the methodology established in Solomon (1987) for generating the time windows. It’s essential to note that customer positions are not sampled from a uniform distribution; instead, they are clustered to replicate real-world scenarios. Further details for instance generation are given in Appendix F.
@@ 4.2 Search performance
We conduct an extensive evaluation of PolyNet’s search performance, benchmarking it against state-of-the-art methods. For the considered routing problems, we train separate models for problem instances of size $100$ and $200$, and then evaluate the models trained on $n$ = $100$ using instances with $100$ and $150$ nodes, and the models trained on $n$ = $200$ using instances with $200$ and $300$ nodes. We can thus assess the model’s capability to generalize to instances that diverge from the training data. Throughout our evaluation, we employ the instance augmentation technique of Kwon et al. (2020).
For the training of PolyNet models, we set the parameter $K$ to $128$ across all problems. We use a learning rate of $10^{-4}$ for the TSP and CVRP and $10^{-5}$ for the CVRPTW. For instances of size nn==$100$, we train our models for $300$ epochs ($200$ for the TSP), with each epoch comprising $4\times 10^{8}$ solution rollouts. Note that we warm-start the training for these models using previously trained POMO models to reduce training times (results for cold-starting training can be found in Appendix C). For instances with nn==$200$, we start training based on the $n$ = $100$ PolyNet models, running $40$ additional training epochs ($20$ for the TSP). To optimize GPU memory utilization, we adjust the batch size separately for each problem and its dimensions.
@@
We use the 10,000 test instances with $n$ = $100$ from Kool et al. (2019) and test sets consisting of $1,000$ instances from Hottung et al. (2021) for $n$ = $150$ and $n$ = $200$. For $n$ = $300$, we generate new instances. As a baseline, we use the exact solver Concorde (Applegate et al., 2006) and the heuristic solver LKH3 (Helsgaun, 2017). Additionally, we also compare to DPDP (Kool et al., 2022b), COMPASS (Chalumeau et al., 2023), and the diversity-focused method MDAM (Xin et al., 2021) with a beam search width of 256.
@@
Method Test (1010K instances) Test (11K instances) Generalization (11K instances)
@@
LKH33 7.765 0.000% 8h 10.687 0.000% 3h 9.346 0.000% 99m 12.949 0.000% 5h
Unguided
POMO  greedy 7.775 0.135% 1m 10.770 0.780% <<1m 9.393 0.494% <<1m 13.216 2.061% 1m
@@
Method Test (1010K instances) Test (11K instances) Generalization (11K instances)
@@
LKH33 15.646 0.53% 6d 22.003 1.09% 25h 19.222 0.88% 20h 28.157 1.51% 34h
Unguided
POMO  greedy 15.754 1.23% 1m 22.194 1.97% <<1m 19.684 3.30% <<1m 28.627 3.21% 1m
@@
Method Test (1010K instances) Test (11K instances) Generalization (11K instances)
@@
Unguided
POMO greedy 13,120 4.67% 1m 19,656 6.70% 1m 18,670 7.25% <<1m 28,022 8.90% 2m
@@
CVRP
Similar to the TSP, we use the test sets from Kool et al. (2019) and Hottung et al. (2021). As a baseline, we use LKH3 (Helsgaun, 2000) and the state-of-the-art (OR) solver HGS (Vidal et al., 2012; Vidal, 2022). We compare to the same learning methods as for the TSP with the addition of DACT (Ma et al., 2021).
CVRPTW
For the CVRPTW, we use the state-of-the-art CO solver PyVRP (Version 0.5.0) (Wouda et al., 2024) as a baseline stopping the search after 1,000 iterations without improvement. We compare to a POMO implementation that we adjusted to solve the CVRPTW by extending the node features with the time windows and the context information used at each decoding step by the current time point. These models are trained for 50,000 epochs, mirroring the training setup used for the CVRP.
4.3 Diversity
Does PolyNet’s training encourage diversity?
To assess the effectiveness of our diversity mechanism, we conduct short training runs of PolyNet for the TSP, CVRP, and CVRPTW with varying values of $K$. As a baseline we report results for the training of the POMO model. For all runs, we use a batch size of $480$ and a learning rate set at $10^{-4}$. To ensure a stable initial state for training runs, we start all runs from a trained POMO model. For PolyNet, we initialize the PolyNet layer weights to zero, minimizing initial randomness. During training, we perform regular evaluations on a separate validation set of 10 00010\,000 instances, sampling $800$ solutions per instance.
Figure 3: Validation performance during training (log scale).
Does PolyNet generate diverse solutions at test time?
In this experiment we evaluate the diversity of PolyNet at test time. As a baseline we use the POMO approach, which enforces diverse solution generation by forcing the first action. For each test instance, we sample $100$ solutions with both approaches and calculate the average broken pairs distance (Prins, 2009), which is an established method to measure diversity for routing problem solutions (see, e.g., Vidal et al. (2012)). The broken pairs distance $d(A,B)$ for two solutions $A$ and $B$ is the number of edges of $A$ that are not in $B$. We measure the diversity of a set of solutions using the average across the broken pairs distances between all solution pairs of the set. The results shown in Table 5 confirm that PolyNet is able to generate more diverse solutions than POMO at test time. An additional visual comparison of the differences in solution diversity between POMO and PolyNet can be found in Appendix D.
Table 4: Solution diversity measured using avg. broken pairs distance.
Method TSP CVRP CVRPTW
PolyNet 4.621 18.535 19.969
POMO 2.400 16.785 18.162
@@ 4.3 Diversity
Gap Time Gap Time Gap Time
PolyNet    Free first move 0.000% 4m 0.49% 5m 2.73% 5m
               Forced first move 0.006% 4m 0.59% 5m 3.00% 6m
Poppy 0.015% 4m 0.78% 5m 3.47% 5m
Does PolyNet encourage useful diversity?
As discussed in the introduction, solution diversity alone is neither inherently beneficial nor difficult to achieve (e.g., by generating random solutions). What truly matters is that the diversity is useful, meaning it contributes to finding high-quality solutions. To assess the diversity of PolyNet we conduct two separate experiments.
First, we evaluate the importance of each of the $K$ learned strategies at test time by counting how often each strategy finds the best solution across 10 00010\,000 test instances. The results, shown in Figure 4, indicate that all $K$ strategies contribute significantly to performance. Even the least effective strategies identify the best solution on 544, 61, and 54 instances for the TSP, CVRP, and CVRPTW, respectively. Note that for the TSP, multiple strategies often converge to the same best solution, likely because most instances are solved to optimality. However, for the CVRP and CVRPTW, the best solution for an instance is typically found by only one strategy. These results demonstrate that all learned strategies are able to generate high-quality solutions and that all strategies contribute meaningfully to the overall sampling performance of PolyNet.
Second, we evaluate the number of distinct first nodes selected by the $K$ different strategies of PolyNet. This provides insights into the diversity achieved by PolyNet, particularly in comparison to POMO, which enforces diversity by selecting different first actions. As shown in Figure 5, PolyNet selects fewer distinct first nodes than POMO. On average, PolyNet selects 24.5, 3.8, and 7.0 distinct first nodes for the TSP, CVRP, and CVRPTW, respectively. Despite selecting fewer different first nodes, PolyNet is able to generate solutions that are more diverse overall than those generated by POMO, as shown in Table 5. This suggests that PolyNet achieves diversity through means other than enforcing the first action.
Figure 4: Contribution of different search strategies. Strategies are sorted based on their contribution.
Figure 5: Frequency plots for the number of unique first nodes.
4.4 Ablation study: Forcing the first move
PolyNet does not force diverse first construction actions and relies solely on its built-in diversity mechanism to select the first node. To assess this approach, we compare the performance of PolyNet with and without forced first move selection when sampling $64\times 100$ solutions per instance.
Table 5 shows the results for all problems with $n$ = $100$. Remarkably, across all scenarios, allowing PolyNet to select the first move yields superior performance compared to forcing the first move. Furthermore, PolyNet with forced first move selection outperforms Poppy (which also enforces the first move), underscoring that PolyNet’s single-decoder architecture delivers better results.
@@ 5 Conclusion
Regarding our approach’s limitations, we acknowledge that the computational complexity of the attention mechanism we employ restricts its applicability to instances with less than $1000$ nodes. However, it is essential to emphasize that the problem sizes examined in this paper for the CVRP(TW) remain challenging for traditional CO solvers and are highly significant in real-world applications. Furthermore, we note that the black-box nature of the PolyNet’s decision-making may be unacceptable in certain decision contexts.
Acknowledgements
André Hottung was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. 521243122. Additionally, we gratefully acknowledge the funding of this project by computing time provided by the Paderborn Center for Parallel Computing (PC2). Furthermore, some computational experiments in this work have been performed using the Bielefeld GPU Cluster. We thank the HPC.NRW team for their support.
@@ Appendix A Experimental results for the FFSP
To train the PolyNet models, we use a learning rate set to $10^{-4}$, and batch sizes of $256$, $128$, and $32$ for instances comprising $n=20$, $50$, and $100$ jobs, respectively. We set the parameter $K$ to $24$ across all problem sizes. The training runs start from trained MatNet models and continue for $100$ (2020 for $n=100$) epochs, with each epoch comprising $24\times 10^{5}$ solution rollouts.
@@
For the FFSP, we use the randomly generated test sets from Kwon et al. (2021) and comparison algorithms. In addition to MatNet (Kwon et al., 2021), we compare our method to CPLEX (CPLEX-Optimization-Studio, 2020) with mixed-integer programming models and metaheuristic solvers. Table 6 provides the results for the FFSP. The gaps are reported with respect to PolyNet with $\times 128$ augmentation result, as optimal solutions are not available. The results demonstrate PolyNet’s consistent improvement in performance across all problem sizes. Notably, PolyNet outperforms MatNet-based solvers for the FFSP on all problem sizes and produces solutions of even higher quality within similar run times. Given the different architecture of MatNet compared to those used for routing problems, the results highlight that the performance gains from PolyNet are not confined to a specific network architecture. Moreover, these gains extend beyond routing problems, indicating PolyNet’s broad applicability across various CO problems.
@@ Appendix B Ablation study: PolyNet layers
Appendix D Solution uniqueness vs. costs plots
In this experiment, we compare the solutions generated by POMO to those generated by PolyNet with respect to their cost and their uniqueness on an instance-by-instance basis. More specifically, for a subset of test instances, we sample $100$ solutions per instance with both approaches and compare the solutions found on the basis of their uniqueness and their cost. We calculate the uniqueness of a solution by using the average of the broken pairs distance (Prins, 2009) to all other 99 generated solutions.
Figures 8-10 shows the results for the first six TSP, CVRP, and CVRPTW test instances. PolyNet generates more unique solutions on most instances. Especially on the TSP the solutions from PolyNet are much more diverse, however this diversity occasionally comes at the expense of solution quality. Since we ultimately only care about the solution with the minimal costs in the generated solution set and, not the average solution quality, this is not a limitation of the method. PolyNet is only able to increase its performance in the minimum cost case by sometimes trying out solutions that a greedier method, like POMO, ignores.
@@
Figure 9: Solution diversity vs. costs for the CVRP
@@
Figure 10: Solution diversity vs. costs for the CVRPTW
@@ Appendix F CVRPTW instance generation
Customer positions are generated using the clustered setting (configuration 2) and customer demands are based on the small, large variance setting (configuration 2). The depot is always centered (configuration 2). It is worth noting that the instance generator samples customer positions within the 2D space defined by $0,999]^{2}$. Independently from the instance generator, vehicle capacities are set at $50$ for instances involving fewer than $200$ customers and increased to $70$ for instances with $200$ or more customers.
To generate the time windows, we adhere to the procedure outlined by Solomon (1987) for instances with randomly clustered customers (i.e., we do not utilize the 3-opt technique to create reference routes). We randomly generate time windows $(e_{i},l_{i})$ for all customers, and set $2400$ as the latest possible time for a vehicle to return to the depot. The time window generation process, as described in Solomon (1987), limits the time windows to ensure feasibility (e.g., by selecting $l_{i}$ so that there is always sufficient time for servicing the customer and returning to the depot). The center of the time window is uniformly sampled from range defined by these limits. We set the maximum width of the time window to $500$ and the service duration to $50$. These parameter values have been deliberately chosen to strike a balance between the constraints of vehicle capacity and time windows, requiring both aspects to be considered during the solution generation proces.